At the beginning of June, Kimber Toy Company budgeted 24,000 toy action figures to be manufactured in June at standard direct materials and direct labor costs as follows:

Direct materials $18,000
Direct labor 14,400
Total $32,400

The standard materials price is $0.50 per pound. The standard direct labor rate is $15.00 per hour. At the end of June, the actual direct materials and direct labor costs were as follows:

Actual direct materials $16,700
Actual direct labor 13,300
Total $30,000

There were no direct materials price or direct labor rate variances for June. In addition, assume no changes in the direct materials inventory balances in June. Kimber Toy Company actually produced 21,600 units during June.

Determine the direct materials quantity and direct labor time variances. Round your per unit computations to two decimal places, if required. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Direct materials quantity variance $
Direct labor time variance $
Answer #1

Direct materials quantity variance = (AQ- SQ)Standard Price

=(33,400 - 32,400)0.50

= 500 (Unfavorable)

Direct labor time variance = (AH - SH) SR

=(886.67 - 864)15

= 340 (Unfavorable)

Working Note

SQ= Standard Quantity = 18,000/0.50 /24,000 x 21,600 = 32,400 Pound

SH = Standard Hours = 14,400/15/24,000 x 21,600 = 864 Hours